# IGNITE: Individualized GeNeration of Imputations in Time-series Electronic health records

### v1

@@ 1 Introduction
Personalized medicine is emerging as a new direction of research, where patients are holistically examined as “individuals” rather than using general population-based symptoms or diagnoses to determine the optimal treatment (Branco and Choupina,, 2021; Lourida and Louridas,, 2021). Patients differ in physiology and response to treatment, where each individual has unique characteristics and patterns that reflect their genetic, molecular and cellular makeup, as well as other environmental and behavioural factors (Lourida and Louridas,, 2021). Electronic health records (EHRs) have opened the door for using real-world patient longitudinal data for a wide range of research such as creating clinical decision support systems for different medical applications  (Cowie et al.,, 2017; Kim et al.,, 2019). EHR adoption has increased worldwide, including in low-middle-income countries (LMIC) (Syzdykova et al.,, 2017). Furthermore, the wealth of data collected in EHRs presents a great source for developing approaches in medicine that tailor the interventions to best fit each individual patient (Abul-Husn and Kenny,, 2019). Moreover, recent works investigated the use of EHRs to generate digital twins of patients, where real-time monitoring, diagnosis, prognosis, and treatment optimization are applied to personalize medicine (Kamel Boulos and Zhang,, 2021; Venkatesh et al.,, 2022).
@@
Data-driven approaches, specifically machine learning models (Xiao et al.,, 2018; Shamout et al.,, 2020), have shown promising results in various clinical applications, especially when using longitudinal EHR data. However, EHRs tend to be multivariate, highly missing and irregularly sampled (Madden et al.,, 2016). Therefore, the success of machine-learning models for personalized medicine highly depends on how the EHR data is represented conditioning on the time-varying physiology, treatments, and the missing values in the data. The existing body of work on personalized medicine mostly emphasizes the utilization of observable data Chen et al., (2019); Wang et al., (2016). However, this approach tends to overlook the crucial aspect of missingness, which is essential for achieving true individualization and obtaining meaningful insights. Missingness in the healthcare domain can occur due to various reasons such as recording errors and device or system failure, irregular sampling and inconsistent medical visits (Kreindler and Lumsden,, 2006). Even high-cost and dangerous to acquire data such as invasive or radiology procedures might be missing due to the risk or lack of feasibility of collection for all patients (Bulas and Egloff,, 2013; Kim et al.,, 2017). Furthermore, the measurement frequency of physiological data could also be related to factors such as patient severity and deterioration (Agniel et al.,, 2018; Ghosheh et al.,, 2022), lack of medical need (Duff et al.,, 2019), bias and low quality of care (Weber et al.,, 2021; Wells et al.,, 2013). The aforementioned characteristics exhibit variability between patients, underscoring the importance of representing the individualized missingness in personalized models.
@@
Handling missingness in longitudinal EHRs is a critical step before conducting statistical or machine learning analysis, as missingness can reach as high as 80% or more in many EHR datasets (Silva et al.,, 2012; Johnson et al.,, 2016; Piri,, 2020). Processing missingness usually involves either data deletion or imputation. Simply deleting recordings or variables with missing values often leads to compromised results and is not a recommended practice, especially when the data is highly missing (Scheffer,, 2002; King et al.,, 1998; Newman and Cottrell,, 2014). Missing value imputation and replacing unobserved values with substitute values methods gained more attention with the advancements in statistical and machine-learning-based models. Current missing-value imputation methods can be classified into three main categories: The first category of methods, referred to as simple imputations, includes replacing the missing values with population mean, median, or personalized Last (observed) Value Carried Forward (LOCF) (Acuna and Rodriguez,, 2004; Shao and Zhong,, 2003). While useful and commonly used in various clinical applications such as clinical trials, imputing by the LOCF presents an issue in observational EHRs, particularly when the feature is never observed for the patient, hence no value to carry forward. The second type of imputation method is a simulation-based statistical class of methods, often called multiple imputation methods. For multiple imputation methods, $N$ complete datasets are simulated from the incomplete dataset by imputing the missing values $N$ times. These completed $N$ datasets are analyzed and then pooled into a single dataset (Jakobsen et al.,, 2017). A common multiple imputation method is multivariate imputation by chained equations (MICE), which is often used by many epidemiological and clinical research (Sterne et al.,, 2009). Despite the usefulness of many approaches, they often fail to handle missingness in longitudinal EHRs as they cannot account for time-series sequential data. The third and most advanced category of imputation methods is based on machine learning algorithms. Matrix Factorization (MF) (Ranjbar et al.,, 2015), Expectation Maximization (EM) (Gold and Bentler,, 2000), K-Nearest Neighbor (KNN) -based imputation (Malarvizhi and Thanamani,, 2012), along with Recurrent Neural Network (RNN) (Cao et al.,, 2018).
@@
Recently, deep generative neural network-based models such as Generative Adversarial Networks (GAN) (Goodfellow et al.,, 2014) and Variational Autoencoder (VAEs) (Kingma and Welling,, 2013) were proposed to generate synthetic data by learning underlying distribution and dynamics of real datasets. The applications of deep generative models in healthcare range from generating synthetic patient records (Choi et al.,, 2017; Li et al.,, 2023), to estimating treatment effects (Yoon et al., 2018b, ), detecting un-diagnosed patients in large-scale  (Li et al.,, 2018), and generating missing value imputations (Yoon et al., 2018a, ; Nazabal et al.,, 2020; Fortuin et al.,, 2020; Luo et al.,, 2019). The generative capabilities of deep generative models make them naturally suitable to generate not only new synthetic records but also missing data imputations, where they show superior performance to most of the commonly used statistical methods in tabular and longitudinal data (Kim et al.,, 2020; Ghosheh et al.,, 2022). Despite their high reported performance, most of the deep generative methods assume missingness is missing at random, which limits their potential in real healthcare applications, where missingness can be informative and reflect individual underlying patient state (Scheffer,, 2002). Furthermore, many of these models can only generate population-level rather than personalized-level information conditioned on the already observed patient data, characteristics and treatments. A comparison of related works is demonstrated in Table 1.
@@
The utilization of binary missingness masks indicating the presence or absence of observations as an extra input feature to prediction models has shown improved performance when compared to models built on observed data only (Sharafoddini et al.,, 2019; Che et al.,, 2018). Despite the utility and predictiveness of the binary missingness masks, they are simplistic and lack indications of the patterns and frequencies of individual-level missingness. We believe in order to generate missing values at a personalized level, robust models that leverage heterogeneous and multiple data types (such as continuous physiological data, discrete treatment data, and individualized missingness masks) are required. Utilizing these mixed data types would capture the non-linear dependencies across high-dimensional longitudinal patient records and generate EHR information that is truly personalized.
To this end, we propose a novel end-to-end generative model for (Individualized GeNeration of Imputations in Time-series Electronic health records (denoted as IGNITE). IGNITE generates a complete patient record by accounting for differences in physiological data, treatment strategies as well as individualized missingness patterns. Adopting a personalized approach to missingness generation, IGNITE could pave the way for creating digital twins for precision medicine, allowing for real-time health monitoring, risk phenotyping and prediction, and treatment strategies at an individual level. We view that deep generative models present a great opportunity not only for imputing values missing at random but also for generating new features that were never measured for the patient. For this purpose, we bring new insights into missingness as a generative task for generating samples and even complete features for patients without any observations for the respective features, which we refer to as feature-wise missingness. An overview of our model is shown in Figure 1. Our main contributions are as follows:
@@
    Individualized Time-series Imputation Model: We propose a generative model comprising dual-variational autoencoders that can impute any personalized missing values in time-series EHR conditioned on discrete treatment (intervention) data and patient characteristics. Our model also includes dual-stage attention networks to improve representation learning across both the feature dimension and temporal dimension.
@@
Figure 1: An overview of the architecture and applications of our proposed model, IGNITE, for generating individualized time-series EHRs. By utilizing treatment data and individualized missingness patterns, IGNITE is capable of generating EHRs that facilitate various applications for personalized medicine.
@@ 1.1 Downstream Task
To evaluate the performance of different imputation models, we used the imputed data in a downstream task — mortality prediction, on each of the three datasets: PhysioNet 2012, HiRID and eICU. The higher the model’s predictive performance, the better the imputation of the data. In the mortality prediction task for the PhysioNet 2012 dataset, our IGNITE performed the best with an area under the receiving operating curve (AUROC) of 0.835, followed by SAITS and transformer AUROC of 0.832 and 0.828, respectively. Models trained on the LOCF-imputed method ranked the worst with an AUROC of 0.758. The various baseline methods were also evaluated in terms of area under the precision-recall curve (AUPRC), which represents discrimination with respect to class imbalance. Similarly, IGNITE achieved the highest AUPRC with a performance of 0.495, followed by Transformer (0.482) and SAITS (0.470), respectively. The results were also consistent on eICU and HiRID datasets, where IGNITE consistently outperformed all the other benchmarks. For instance, IGNITE achieved an AUROC of 0.774 in eICU and 0.968 in HiRID, respectively. A similar trend was observed in AUPRC results for both datasets. It is worth noting that the eICU imputations generated by IGNITE significantly improved the performance with an increase of over 10% and 20% when compared to the next best or the worst models, respectively.
Table 2: Performance for a mortality prediction task using an LSTM model reported in terms of AUROC and AURPC
@@
LOCF 0.758 0.595 0.901 0.360 0.205 0.696
MICE 0.818 0.551 0.897 0.464 0.167 0.661
GP-VAE 0.767 0.724 0.963 0.382 0.275 0.805
Transformer 0.828 0.758 0.945 0.482 0.314 0.815
BRITS 0.819 0.731 0.953 0.461 0.232 0.888
SAITS 0.832 0.675 0.935 0.470 0.208 0.878
IGNITE 0.835 0.774 0.968 0.495 0.389 0.892
The baseline imputation methods were also evaluated in terms of performance with varied percentages and types of missingness for feature-wise and sample-wise missingness, respectively. The three datasets were stratified based on the percentage of overall feature-wise missingness, where $z\%$ of features were never observed. As shown in Table 3, we note that for the PhysioNet 2012 dataset, only $\leq$ 25% and 25% - 75% missingness were evaluated as not enough training samples were available with $\geq$ 75% missingness. IGNITE achieved the highest performance with an AUROC of 0.790, AUPRC of 0.561 for $\leq$ 25%, and 0.791 and 0.346 in terms of AUROC and AUPRC in the 25% - 75% task, respectively. Similarly, IGNITE consistently outperformed the other baselines regarding AUPRC for eICU in the population subset with $\leq$ 25% of feature-wise missingness. As for the population with feature-wise missingness of 25%- 75% for eICU, the highest performance was achieved by SAITS, followed by IGNITE. For HiRID, on the other hand, the highest performance of the population with feature-wise missingness of 25%- 75% was achieved by the model trained on the data imputed by the Transformer model in terms of AUROC, in terms of AURPC, the highest performance was achieved by IGNITE. As most of the population has < 75% feature-wise missingness, HiRID and PhysioNet 2012 were not used for the experiments with > 75% missingness. For eICU, IGNITE achieved the highest performance, followed by SAITS in terms of AUROC for the population with > 75% missingness. We note that the performance of GP-VAE, MICE, and LOCF imputation ranked last across multiple percentages of missingness, indicating the lack of robustness.
@@
To evaluate the population for varying sample-wise missingness, where values are not observed across time and features, we stratified the dataset into three groups $\leq$ 25%, 25%- 75%, and $\geq$ 75% sample-wise missingness. However, in all three datasets, no patients had $\leq$ 25% sample-wise missingness, indicating the high missingness in EHRs. The high prevalence of missingness in all the included datasets resulted in two main subgroups for 25%- 75%, and $\geq$ 75% sample-wise missingness for PhysioNet 2012 and HiRID. On the other hand, more than 95% of the samples in the eICU dataset had more than 75% sample-wise missingness, resulting in one subgroup for eICU. In terms of performance, IGNITE consistently outperformed all the baselines in terms of AUROC in all three datasets. Precisely, IGNITE achieved the highest performance of 0.816, followed by BRITS for the 25%- 75% group in the PhysioNet 2012 dataset. For AURPC, IGNITE achieved the second-highest performance with 0.419, where the highest performance was achieved by GP-VAE. Similar trends were observed in the eICU and HiRID datasets, with an AUROC reaching 0.959 for the HiRID dataset for samples with > 75% sample-wise missingness. While SAITS and Transformer showed high performance across all population subsets, the performance deteriorated for the population with low sample-wise missingness. We also observe that the outcome of mortality was less prevalent in patients with higher overall sample-wise missingness, 13.4% when compared to those with lower sample-wise missingness 25.1% in the PhysioNet 2012 dataset. Similar trends were also observed in the HiRID dataset, where the outcome prevalence decreased from 23.9% to 5.43% for the samples with high sample-wise missingness.

### v2

@@ 1 Introduction
Personalized medicine is emerging as a new direction of research, in which patients are holistically examined as “individuals” rather than using symptoms or diagnoses based on the general population to determine the optimal treatment (Branco and Choupina,, 2021; Lourida and Louridas,, 2021). Patients differ in physiology and response to treatment, where each individual has unique characteristics and patterns that reflect their genetic, molecular, and cellular makeup, as well as other environmental and behavioral factors (Lourida and Louridas,, 2021). Electronic health records (EHRs) have opened the door to the use of longitudinal data from real-world patients for a wide range of research, such as creating clinical decision support systems for different medical applications  (Cowie et al.,, 2017; Kim et al.,, 2019). The adoption of EHR has increased worldwide, including in low-middle-income countries (LMICs) (Syzdykova et al.,, 2017). Furthermore, the wealth of data collected in EHRs presents a great source for developing approaches in medicine that tailor interventions to best suit each individual patient (Abul-Husn and Kenny,, 2019). In addition, recent work investigated the use of EHRs to generate digital twins of patients, where real-time monitoring, diagnosis, prognosis, and treatment optimization are applied to personalize the medicine (Kamel Boulos and Zhang,, 2021; Venkatesh et al.,, 2022).
@@
Data-driven approaches, specifically machine learning models (Xiao et al.,, 2018; Shamout et al.,, 2020), have shown promising results in various clinical applications, especially when using longitudinal EHR data. However, EHRs tend to be multivariate, highly missing, and irregularly sampled (Madden et al.,, 2016). Therefore, the success of machine-learning models for personalized medicine is highly dependent on how the EHR data are represented conditioning on the time-varying physiology, treatments, and the missing values in the data. The existing body of work on personalized medicine mostly emphasizes the utilization of observable data Chen et al., (2019); Wang et al., (2016). However, this approach tends to overlook the crucial aspect of missingness, which is essential to achieve true individualization and to obtain meaningful insights. Missingness in the healthcare domain can occur due to various reasons, such as recording errors and device or system failure, irregular sampling, and inconsistent medical visits (Kreindler and Lumsden,, 2006). Even high-cost and dangerous data acquisition, such as invasive or radiological procedures, may be missing due to the risk or lack of feasibility of collection for all patients (Bulas and Egloff,, 2013; Kim et al.,, 2017). Furthermore, the measurement frequency of physiological data could also be related to factors such as patient severity and deterioration (Agniel et al.,, 2018; Ghosheh et al.,, 2022), lack of medical need (Duff et al.,, 2019), bias and low quality of care (Weber et al.,, 2021; Wells et al.,, 2013). The aforementioned characteristics exhibit variability between patients, underscoring the importance of representing the individualized missingness in personalized models.
@@
The handling of missingness in longitudinal EHRs is a critical step before performing statistical or machine learning analysis, as missingness can reach as high as 80% or more in many EHR datasets (Silva et al.,, 2012; Johnson et al.,, 2016; Piri,, 2020). Processing missingness usually involves data deletion or imputation. Simply deleting recordings or variables with missing values often leads to compromised results and is not recommended practice, especially when the data is highly missing (Scheffer,, 2002; King et al.,, 1998; Newman and Cottrell,, 2014). Missing value imputation and replacing unobserved values with substitute values methods gained more attention with the advancements in statistical and machine-learning-based models. Current missing-value imputation methods can be classified into three main categories: The first category of methods, known as simple imputations, includes replacing missing values with population mean, median, or personalized Last (observed) Value Carried Forward (LOCF) (Acuna and Rodriguez,, 2004; Shao and Zhong,, 2003). Although useful and commonly used in various clinical applications such as clinical trials, imputing by the LOCF presents an issue in observational EHRs, particularly when the feature is never observed for the patient, hence no value to carry forward. The second type of imputation method is a simulation-based statistical class of methods, often called multiple imputation methods. For multiple imputation methods, $N$ complete datasets are simulated from the incomplete dataset by imputing the missing values $N$ times. These completed $N$ datasets are analyzed and then pooled into a single dataset (Jakobsen et al.,, 2017). A common multiple imputation method is multivariate imputation by chained equations (MICE), which is often used by many epidemiological and clinical research (Sterne et al.,, 2009). Despite the usefulness of many approaches, they often fail to handle missingness in longitudinal EHRs as they cannot account for time-series sequential data. Although MICE is useful for certain imputation tasks, they may not be as effective in modeling the sequential nature of time-series data as they read the data in a flattened form. This approach often treats each time point independently, which can lead to suboptimal imputations when dealing with the continuous and interdependent nature of time-series data. The third and most advanced category of imputation methods is based on machine learning algorithms. Matrix Factorization (MF) (Ranjbar et al.,, 2015), Expectation Maximization (EM) (Gold and Bentler,, 2000), K-Nearest Neighbor (KNN) -based imputation (Malarvizhi and Thanamani,, 2012), along with Recurrent Neural Network (RNN) (Cao et al.,, 2018).
@@
Recently, deep generative neural network-based models such as Generative Adversarial Networks (GAN) (Goodfellow et al.,, 2014) and Variational Autoencoder (VAEs) (Kingma and Welling,, 2013) were proposed to generate synthetic data by learning the underlying distribution and dynamics of real datasets. The applications of deep generative models in healthcare range from generating synthetic patient records (Choi et al.,, 2017; Li et al.,, 2023), to estimating treatment effects (Yoon et al., 2018b, ), detecting un-diagnosed patients in large-scale  (Li et al.,, 2018), and generating missing value imputations (Yoon et al., 2018a, ; Nazabal et al.,, 2020; Fortuin et al.,, 2020; Luo et al.,, 2019). The generative capabilities of deep generative models make them naturally suitable to generate not only new synthetic records but also missing data imputations, where they show superior performance to most of the commonly used statistical methods in tabular and longitudinal data (Kim et al.,, 2020; Ghosheh et al.,, 2024; Wang et al.,, 2024). Despite their high reported performance, most deep generative methods assume missingness is missing at random, which limits their potential in real healthcare applications, where missingness can be informative and reflect individual underlying patient state (Scheffer,, 2002). Furthermore, many of these models can only generate population-level rather than personalized-level information conditioned on the already observed patient data, characteristics and treatments.
In our evaluation of IGNITE, we have focused on multivariate time-series imputation within the context of electronic health records (EHRs), where continuous monitoring and recording of patient information over time is prevalent. This necessitates the inclusion of methods that are directly comparable in this domain. Although influential methods such as MD-MTS (Xu 2020) and SMILES (Zhang 2020) have been cited in the literature, these methods are primarily focused on static data imputation or specific to non-multivariate time-series data. Our work specifically targets the unique challenges of multivariate time-series imputation in EHRs. A comparison of previous works is shown in Table 1.
@@
The utilization of binary missingness masks that indicate the presence or absence of observations as an additional input feature to prediction models has shown improved performance compared to models built on observed data only (Sharafoddini et al.,, 2019; Che et al.,, 2018). Despite the utility and predictiveness of binary missingness masks, they are simplistic and lack indications of the patterns and frequencies of individual-level missingness. We believe that to generate missing values at a personalized level, robust models that leverage heterogeneous and multiple data types (such as continuous physiological data, discrete treatment data, and individualized missingness masks) are required. Utilizing these mixed data types would capture the non-linear dependencies across high-dimensional longitudinal patient records and generate EHR information that is truly personalized.
To this end, we propose a novel end-to-end generative model for (Individualized GeNeration of Imputations in Time-series Electronic health records (denoted as IGNITE). IGNITE generates a complete patient record taking into account differences in physiological data, treatment strategies as well as individualized missingness patterns. Adopting a personalized approach to missingness generation, IGNITE could pave the way for creating digital twins for precision medicine, allowing real-time health monitoring, risk phenotyping and prediction, and treatment strategies at an individual level. We view that deep generative models present a great opportunity not only for imputing values missing at random but also for generating new features that were never measured for the patient. For this purpose, we bring new insights into missingness as a generative task for generating samples and even complete features for patients without any observations of the respective features, which we refer to as feature-wise missingness. By combining these elements, IGNITE not only addresses the persistent challenges of incompleteness and variability of data in EHRs, but also provides more accurate and personalized risk assessment and treatment strategies, which contributes significantly to the improvement of patient outcomes and the optimization of clinical workflows. An overview of our model is shown in Figure 1. Our main contributions are as follows:
@@
    Individualized Time-series Imputation Model: We propose a generative model comprising dual-variational autoencoders that can impute any personalized missing values in time-series EHR conditioned on discrete treatment (intervention) data and patient characteristics. Our model also includes dual-stage attention networks to improve representation learning across both the feature dimension and the temporal dimension.
@@
Figure 1: An overview of the architecture and applications of our proposed model, IGNITE, for generating individualized time-series EHRs. The evidence lower bound (ELBO) is calculated for the observed value only for the upper VAE and for the augmented data from the full individualized missingness mask (IMM) in the lower VAE. By utilizing treatment data and individualized missingness patterns, IGNITE is capable of generating EHRs that facilitate various applications of personalized medicine.
@@ 1.1 Downstream Task
To evaluate the performance of different imputation models, we used the imputed data in a downstream task — mortality prediction, on each of the three datasets: PhysioNet 2012, HiRID and eICU. In all models, we first imputed the data and then used the imputed data for evaluation in the respective downstream task. The higher the predictive performance of the model, the better the imputation of the data. In the mortality prediction task for the PhysioNet 2012 dataset, our IGNITE performed the best with an area under the receiving operating curve (AUROC) of 0.834, followed by TimesNet and transformer AUROC of 0.831 and 0.827, respectively. Models trained on the LOCF-imputed method were ranked the worst with an AUROC of 0.777. The various baseline methods were also evaluated in terms of area under the precision-recall curve (AUPRC), which represents discrimination with respect to class imbalance. The results were also consistent on the eICU and HiRID datasets, where IGNITE consistently outperformed all other benchmarks. For instance, IGNITE achieved an AUROC of 0.822 in eICU and 0.968 in HiRID, respectively. A similar trend was observed in the AUPRC results for both datasets. It is worth noting that the eICU imputations generated by IGNITE significantly improved the performance with an increase of more than 10% and 20% when compared to the next best or worst models, respectively. We also implemented Wilcoxon Signed-Ranked test to assess for the statistical significance and show the results .
Table 2: Performance for a mortality prediction task using an LSTM model reported in terms of AUROC and AURPC. $\ast$ indicates that IGNITE is statistically significant ($p<0.05$) when compared to the corresponding benchmark.
@@
LOCF 0.777 (0.014)* 0.736 (0.001)* 0.935 (0.007)* 0.389 (0.025)* 0.316 (0.015)* 0.593 (0.031)*
MICE 0.823 (0.009)* 0.765 (0.014)* 0.939 (0.007)* 0.469 (0.025)* 0.289 (0.034)* 0.563 (0.023)*
GP-VAE 0.780 (0.013)* 0.798 (0.009)* 0.948 (0.006)* 0.371 (0.024)* 0.397 (0.036)* 0.583 (0.025)*
Transformer 0.831 (0.010)* 0.729 (0.008)* 0.961 (0.004)* 0.477 (0.028)* 0.249 (0.017)* 0.658 (0.022)*
BRITS 0.818 (0.010)* 0.735 (0.001)* 0.959 (0.005)* 0.452 (0.025)* 0.318 (0.015)* 0.641 (0.027)*
TimesNet 0.827 (0.001)* 0.750 (0.010)* 0.956 (0.004)* 0.352 (0.014)* 0.305 (0.014)* 0.640 (0.022)*
SAITS 0.821 (0.010)* 0.608 (0.007)* 0.961 (0.005)* 0.467 (0.035)* 0.115 (0.008)* 0.655 (0.030)*
IGNITE 0.834 (0.009) 0.822 (0.008) 0.968 (0.004) 0.458 (0.032) 0.345 (0.014) 0.723 (0.030)
The baseline imputation methods were also evaluated in terms of performance with varied percentages and types of missingness for feature-wise and sample-wise missingness, respectively. The three datasets were stratified based on the percentage of overall feature-wise missingness, where $z\%$ of features were never observed. As shown in Table 3, we note that for the PhysioNet 2012 dataset, only $\leq$ 25% and 25% - 75% missingness were evaluated, as there were not enough training samples with $\geq$ 75% missingness. IGNITE achieved the highest performance with an AUROC of 0.790, an AUPRC of 0.561 for $\leq$ 25%, and 0.791 and 0.346 in terms of AUROC and AUPRC in the 25% - 75% task, respectively. Similarly, IGNITE consistently outperformed the other baselines in assessing AUPRC for eICU in the population subset with $\leq$ 25% of feature-wise missingness. For the population with feature-wise missingness of 25%- 75% for eICU, SAITS achieved the highest performance, followed by IGNITE. For HiRID, on the other hand, the highest performance of the population with feature-wise missingness of 25%- 75% was achieved by the model trained on the data imputed by the Transformer model in terms of AUROC, in terms of AURPC, IGNITE achieved the highest performance. As most of the population has < 75% feature-wise missingness, HiRID and PhysioNet 2012 were not used for the experiments with > 75% missingness. For eICU, IGNITE achieved the highest performance, followed by SAITS in terms of AUROC for the population with > 75% missingness. We note that the performance of the GP-VAE, MICE, and LOCF imputation ranked last across multiple percentages of missingness, indicating the lack of robustness.
@@
To evaluate the population for varying sample-wise missingness, where values are not observed across time and features, we stratified the dataset into three groups $\leq$ 25%, 25%- 75%, and $\geq$ 75% sample-wise missingness. However, in all three datasets, no patients had $\leq$ 25% sample-wise missingness, indicating the high missingness in EHRs. The high prevalence of missingness in all the included datasets resulted in two main subgroups for 25%- 75%, and $\geq$ 75% sample-wise missingness for PhysioNet 2012 and HiRID. On the other hand, more than 95% of the samples in the eICU dataset had more than 75% sample-wise missingness, resulting in one subgroup for eICU. In terms of performance, IGNITE consistently outperformed all baselines in terms of AUROC in all three datasets. Specifically, IGNITE achieved the highest performance of 0.816, followed by BRITS for the 25%- 75% group in the PhysioNet 2012 dataset. For AURPC, IGNITE achieved the second-highest performance with 0.419, where the highest performance was achieved by GP-VAE. Similar trends were observed in the eICU and HiRID datasets, with an AUROC reaching 0.959 for the HiRID dataset for samples with > 75% sample-wise missingness. Although SAITS and Transformer showed high performance across all population subsets, the performance deteriorated for the population with low sample-wise missingness. We also observed that mortality outcome was less prevalent in patients with higher overall sample-wise missingness, 13.4% when compared to those with lower sample-wise missingness 25.1% in the PhysioNet 2012 dataset. Similar trends were also observed in the HiRID dataset, where the outcome prevalence decreased from 23.9% to 5.43% for the samples with high sample-wise missingness.